Circular No. 7702

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001eb IN NGC 1589 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles

(Coddenham, Suffolk, England) of an apparent supernova (mag 16.3) on a single unfiltered CCD image taken (limiting mag ∼ 18.0) on Aug. 29.132 UT in the course of the U.K. Nova/Supernova Patrol with a 0.36-m Schmidt-Cassegrain telescope. SN 2001eb is located at α = 4^{h}30^{m}45^{s}.10, δ = +0^{o}52′17′′.9 (equinox 2000.0), which is 7′′ west and 28′′ north of the center of NGC 1589. The new object is embedded in the spiral arms and therefore appears slightly diffuse. In response to an appeal from ‘The Astronomer’ group, D. Balam, (Victoria, BC, Canada) imaged the suspect on Aug. 30.509 with the 1.82-m Plaskett telescope, reporting that the brightness appeared similar to that found by Boles. Boles adds that the object is missing from his master image of 1999 Feb. 17 (limiting mag 17.5) and from second Palomar Sky Survey images taken in 1990 (red, limiting mag 19.5) and 1991 (blue, limiting mag 19.5). Hurst also notes that SN 2001eb is not present on an image of this galaxy shown in J. C. Vickers’ Deep Space CCD Atlas: North (1993, p. 43; limiting mag ∼ 18).

SUPERNOVA 2001ec IN PGC 74077 D. Weisz and W. D. Li, University of California at Berkeley, reports

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.8) in an unfiltered image taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 31.5 UT. The new object is confirmed at mag ∼ 18.3 in an earlier image taken on Aug. 19.5. SN 2001ec is located at α = 2^{h}12^{m}00^{s}.43, δ = +31^{o}32′43′′.8 (equinox 2000.0), which is 5′′.1 east and 5′′.5 north of the nucleus of PGC 74077. A KAIT image obtained on Feb. 4.2 showed nothing at this position (limiting magnitude ∼ 19.0).

V4739 SAGITTARII CCD V magnitudes: Aug. 28.087 UT, 8.66 (J. D. West, Mulvane, KS);

29.085, 9.29 (West); 30.098, 10.24 (West); 31.375, 11.06 (A. C. Gilmore, Mount John Observatory); Sept. 1.412, 11.38 (Gilmore). Additional photometry by Gilmore (cf. IAUC 7698): Aug. 31.375, U − B = −0.65, B − V = −0.36, V −R = +1.53, V −I = +1.15 (air mass 1.04); Sept. 1.412, 11.38, −0.66, −0.39, +1.62, +1.20 (1.08).

2001 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green